# On the Non-Identifiability of Steering Vectors in Large Language Models

### v2

@@ Abstract
Activation steering methods are widely used to control large language model (LLM) behavior and are often interpreted as revealing meaningful internal representations. This interpretation assumes steering directions are identifiable and uniquely recoverable from input–output behavior. We show that, under white-box single-layer access, steering vectors are fundamentally non-identifiable due to large equivalence classes of behaviorally indistinguishable interventions. Empirically, we show that orthogonal perturbations achieve near-equivalent efficacy with negligible effect sizes across multiple models and traits. Critically, we show that the non-identifiability is a robust geometric property that persists across diverse prompt distributions. These findings reveal fundamental interpretability limits and highlight the need for structural constraints beyond behavioral testing to enable reliable alignment interventions. [^1] Code: https://github.com/sohv/non-identifiability
@@
Sohan Venkateshthanks: Correspondence: soh.venkatesh@gmail.com     Ashish Mahendran Kurapath Manipal Institute of Technology Bengaluru
@@ 1 Introduction
Despite growing adoption in interpretability and alignment research, the theoretical foundations of persona steering remain poorly understood. Most existing methods implicitly assume that extracted steering vectors correspond to meaningful, uniquely determined latent factors and that these factors can be directly manipulated to achieve reliable control. However, classical results in latent variable modeling and causal inference show that such assumptions are often unjustified without additional structural constraints (Hyvärinen and Pajunen, 1999; Shimizu et al., 2006; Schölkopf et al., 2021; Locatello et al., 2019).
@@
Figure 1: Geometric illustration of steering vector non-identifiability. Two geometrically distinct vectors, the original steering vector $v$ (left) and a perturbed version $v+v_{\perp}$ where $v_{\perp}\in\mathrm{ker}(J)$ (right), produce identical observable outputs. The null-space component (red, unobservable region) can be arbitrarily modified without changing outputs, demonstrating the existence of large equivalence classes of behaviorally indistinguishable steering directions.
We make three contributions. First, we provide a formal identifiability analysis of persona vector steering, proving that under white-box single-layer access, persona vectors are fundamentally non-identifiable: infinitely many geometrically distinct directions induce identical observable behavior via null-space ambiguity. Second, we empirically demonstrate that extracted steering vectors contain substantial non-identifiable components, with orthogonal perturbations achieving 95–100% of the original steering efficacy across multiple models and traits. Third, we validate through multi-environment experiments that this equivalence persists under distribution shift, confirming non-identifiability is a robust geometric property rather than an artifact of limited prompt diversity. These results clarify when alignment methods can be trusted and which structural assumptions are required for reliable control.
@@ 2 Related Work
Causal and latent variable identifiability. Classical results show that latent variable models are non-identifiable without structural assumptions (Hyvärinen and Pajunen, 1999; Shimizu et al., 2006; Kruskal, 1977). Recent work in causal representation learning extends these ideas to deep learning settings (Schölkopf et al., 2021; Ahuja et al., 2022; Locatello et al., 2019), establishing conditions under which latent factors can be recovered from high-dimensional observations.
Probing and representation learning. Work on probing classifiers (Pimentel et al., 2020; Ravfogel et al., 2020; Elazar et al., 2021) shows that linear directions can succeed for reasons unrelated to target information. The linear representation hypothesis (Park et al., 2023; Elhage et al., 2022) suggests concepts correspond to directions in activation space but lacks identifiability guarantees.
Activation editing in LLMs. Methods such as representation engineering (Zou et al., 2023), contrastive activation addition (Rimsky et al., 2024; Turner et al., 2024) and activation patching (Meng et al., 2022) manipulate model internals but do not address whether control directions are uniquely determined. Work on neural network symmetries (Entezari et al., 2021; Dinh et al., 2017) motivates our null-space analysis.
Robustness and invariant representations. Work on distribution shift and domain adaptation (Rabanser et al., 2019; Lipton et al., 2018) establishes that learned representations often exploit spurious correlations. Methods for learning invariant representations (Arjovsky et al., 2019; Krueger et al., 2021) aim to extract features stable across domains. Recent work on multi-environment causal representation learning (Ahuja et al., 2022; Von Kügelgen et al., 2021; Rojas-Carulla et al., 2018) proves that diverse training environments enable identification of causal factors under structural assumptions.
Mechanistic interpretability. Work on mechanistic interpretability (Olsson et al., 2022; Elhage et al., 2021; Nanda et al., 2023) studies circuits and features in transformers, often assuming the existence of interpretable directions without formalizing identifiability constraints. Our theoretical analysis clarifies when such directions are well-defined.
@@ 3.1 Formal Model
Latent persona variable.
@@ 3.2 Observational Regimes
  • Regime 1: Black-box input–output. The researcher observes only $(x,y)$ pairs, where $y$ is generated text. There is no access to internal representations. This is the weakest regime and corresponds to behavioral evaluation.
  • Regime 2: White-box single-layer access. The researcher can observe or manipulate activations $h_{\ell}(x)$ at a chosen layer $\ell$. This is the standard setting for most steering work and includes extracting vectors from contrastive prompt pairs.
@@ 3.4 Assumptions
We make the following assumptions explicit:
A1 (Smoothness).
The functions $g_{\ell}$ and $F_{\ell\to L}$ are differentiable almost everywhere, enabling local linear approximation via Jacobians $J_{\ell}(x)=\frac{\partial o}{\partial h_{\ell}}(x)$.
A2 (Identifiable prompts).
The prompt distribution $p(x)$ has sufficient variability to probe different aspects of the latent persona $z$. Formally, the support of $p(x)$ is rich enough that different persona values induce distinguishable activation patterns $h_{\ell}(x)$.
A3 (Non-degeneracy).
The Jacobian $J_{\ell}(x)$ has rank at least $k\geq 1$ for typical $x\sim p(x)$, meaning steering can affect outputs. This excludes pathological cases where all perturbations to $h_{\ell}$ are ignored by subsequent layers.
We do not assume:
  • Statistical independence between $z$ and $x$ (confounding is allowed)
  • Linearity of $g_{\ell}$ in $z$
  • Uniqueness of the latent representation without additional structure
These assumptions are mild and satisfied by standard transformer architectures under typical operating conditions.
@@ 4.1 Identifiability
Definition 1 (Parameter Identifiability).
@@
Definition 2 (Observational Equivalence).
Two steering vectors $v$ and $v^{\prime}$ are observationally equivalent in regime $\mathcal{R}$ if they produce identical distributions over all quantities observable in $\mathcal{R}$.
For Regime 2 (white-box single-layer access):
@@
$=F_{\ell\to L}(h_{\ell}(x)+\alpha v^{\prime})\;\forall x,\alpha.$
@@ 4.2 Symmetries and Gauge Freedom
Null space ambiguity.
@@ 5 Main Results
Proposition 1.
Under Assumptions A1–A3, in Regime 2 (white-box single-layer access) without additional structural constraints, persona vectors are not identifiable. Specifically, for any steering vector $v\in\mathbb{R}^{d}$, there exist infinitely many vectors $v^{\prime}\not\propto v$ that are observationally equivalent.
Proof Sketch.
We establish non-identifiability via two complementary arguments: null-space ambiguity and reparameterization symmetry.
Null-space ambiguity (primary mechanism). Under the local linear approximation $o\approx o_{0}+\alpha J_{\ell}v$, any vector $v^{\prime}=v+v_{0}$ where $v_{0}\in\ker(J_{\ell})$ satisfies $J_{\ell}v^{\prime}=J_{\ell}v$, and thus produces identical linearized outputs. This establishes non-identifiability even without invoking global reparameterization symmetries.
Reparameterization symmetry. Neural networks exhibit inherent parameter symmetries arising from overparameterization. Consider an invertible linear transformation $T(h)=Ah$ where $A\in\mathbb{R}^{d\times d}$. For layer $\ell+1$ with weight matrix $W_{\ell+1}$ and bias $b_{\ell+1}$, the computation $h_{\ell+1}=\sigma(W_{\ell+1}h_{\ell}+b_{\ell+1})$ can be equivalently expressed through reparameterized weights $W^{\prime}_{\ell+1}=W_{\ell+1}A^{-1}$:
$h^{\prime}_{\ell+1}$ $=\sigma\!\left(W^{\prime}_{\ell+1}Ah_{\ell}+b_{\ell+1}\right)$ (7)
$=\sigma\!\left(W_{\ell+1}h_{\ell}+b_{\ell+1}\right)$
$=h_{\ell+1}$
Under this reparameterization, steering with $v^{\prime}=Av$ on the transformed representation $h^{\prime}_{\ell}=Ah_{\ell}$ produces identical outputs to steering with $v$ on $h_{\ell}$. For any invertible matrix $A\neq cI$ (not a scalar multiple of identity), we obtain $v^{\prime}\not\propto v$, yet these vectors are observationally equivalent. Since the space of such matrices has dimension $d^{2}-1$, infinitely many distinct reparameterizations exist.
Practical vs. theoretical relevance: The reparameterization argument establishes non-identifiability as a fundamental property of the model class, independent of any specific instantiation. While these symmetries cannot be exploited in frozen models without modifying weights, their existence proves that identifiability fails even in principle. For analyzing deployed models, the null-space mechanism provides direct, operationally relevant non-identifiability.
Corollary 1.1 (Null-space equivalence).
Under the linear approximation $o\approx o_{0}+\alpha J_{\ell}v$, any vector $v^{\prime}=v+v_{0}$ where $v_{0}\in\ker(J_{\ell})$ is observationally equivalent to $v$, since $J_{\ell}v^{\prime}=J_{\ell}v+J_{\ell}v_{0}=J_{\ell}v$.
Remark (Null-space dimensionality).
The null space $\ker(J_{\ell})$ is typically high-dimensional in practice. For a Jacobian $J_{\ell}\in\mathbb{R}^{V\times d}$ with vocabulary size $V$ and hidden dimension $d$, the maximum possible rank is $\min(V,d)$. In modern language models, $V\approx 50{,}000$ and $d\approx 4{,}000$, so the maximal rank is $d$. However, empirical evidence shows that (1) neural network representations exhibit low intrinsic dimensionality (Li et al., 2018; Maennel et al., 2018), and (2) output sensitivity concentrates in a small number of Hessian directions (Ghorbani et al., 2019). These findings suggest the Jacobian $J_{\ell}$ has effective rank substantially smaller than $d$, implying a high-dimensional null space:
$\dim(\ker(J_{\ell}))=d-\mathrm{rank}(J_{\ell})$ (8)
@@ 6 Empirical Validation
We now validate that the non-identifiability characterized in Proposition 1 (Section 5) manifests in contemporary language models. Our empirical experiments test a behavioral consequence of the theoretical prediction: the existence of large equivalence classes of semantically indistinguishable steering directions, without directly estimating the Jacobian null space.
@@ 6.1 Experimental Setup
Models and Layers.
We evaluate two open-weight instruction-tuned language models to test generality across architectures and scales: Qwen2.5-3B-Instruct (layer $\ell=12$) and Llama-3.1-8B-Instruct (layer $\ell=16$). For both models, we focus on middle layers ($\ell=L/2$), following standard practice (Chen et al., 2025; Konen et al., 2024; Sun et al., 2025) in steering research.
Persona traits.
@@
Steering vector extraction.
@@
Semantic probes.
We evaluate semantic equivalence using trait-specific scoring functions $\phi(o)$ that map generated text to real-valued scores. For formality, politeness and humor, we use lexical heuristics based on formal/informal markers, polite/rude markers and humorous/serious markers respectively, combined with sentence length and stylistic features. These return scores in $[0,1]$ where 1 is maximally formal/polite/humorous.
Orthogonality test methodology.
@@
Sample size design.
@@
Scale invariance test.
@@ 6.2 Orthogonal Perturbation Test Results
Table 1: Empirical validation of non-identifiability across models, traits and sample sizes. Cohen’s $d$ measures effect size between $v$ and $v+v_{\perp}$ (lower = more equivalent). All values show negligible differences ($d<0.2$), confirming observational equivalence.
@@
With $n=10$ seeds, the mean Cohen’s $d$ is 0.080 for Qwen2.5-3B and 0.100 for Llama-3.1-8B, both well below the threshold for small effects ($d=0.2$) and firmly in the negligible range. The “Perp-Only Effect” column measures the efficacy of steering with pure orthogonal components ($v_{\perp}$ alone, without $v$) relative to the extracted vector—values near $100\%$ indicate that orthogonal components achieve equivalent behavioral impact.
Statistical stability across sample sizes. The convergence between $n=5$ and $n=10$ results demonstrates robustness across both models. For Qwen2.5-3B, effect sizes change by <0.07<0.07 as sample size increases, with all traits showing tightening confidence intervals: formality ($0.144\to 0.075$), politeness ($0.112\to 0.092$), humor ($0.103\to 0.072$). For Llama-3.1-8B, we observe similar stability: formality ($0.052\to 0.096$), politeness ($0.074\to 0.085$), humor ($0.159\to 0.119$). This consistency demonstrates that the observed equivalence is not a sampling artifact but a stable property of the steering geometry.
Cross-model consistency. The close agreement between Qwen2.5-3B ($d=0.080$) and Llama-3.1-8B ($d=0.100$) demonstrates that observational equivalence is not model-specific. Despite differences in architecture, scale (3B vs. 8B parameters) and hidden dimension ($d=2048$ vs. $d=4096$), both models exhibit nearly identical patterns of non-identifiability. The consistency across traits within each model further confirms that the phenomenon is general rather than an artifact of specific semantic domains or model implementations.
Visualizing orthogonal component equivalence. Figure 2 illustrates the perp-only effect ratios across all traits and models using $n=10$ orthogonal seeds. The tight clustering around the perfect equivalence line (dashed red at $1.0$) demonstrates that pure orthogonal components achieve nearly identical steering efficacy to the extracted vectors.
@@
Figure 2: Perp-only effect ratios ($v_{\perp}$ efficacy / $v$ efficacy) for $n=10$ orthogonal seeds. Values near 1.0 (dashed red line) indicate perfect equivalence.

### v3

@@ Abstract
Activation steering methods are widely used to control large language model (LLM) behavior and are often interpreted as revealing meaningful internal representations. This interpretation assumes that steering directions are identifiable and uniquely recoverable from input–output behavior. We show that, under white-box single-layer access, steering vectors are fundamentally non-identifiable due to large equivalence classes of behaviorally indistinguishable interventions. Empirically, we find that orthogonal perturbations achieve near-equivalent efficacy with negligible effect sizes across multiple models and traits. Critically, we show that non-identifiability is a robust geometric property that persists across diverse prompt distributions. These findings reveal fundamental interpretability limits and highlight the need for structural constraints beyond behavioral testing to enable reliable alignment interventions.
@@
Sohan Venkatesh1thanks: Correspondence: soh.venkatesh@gmail.com        Ashish Mahendran Kurapath1 1Manipal Institute of Technology Bengaluru
@@ 1 Introduction
Despite growing adoption in interpretability and alignment research, the theoretical foundations of persona steering remain poorly understood. Most existing methods implicitly assume that extracted steering vectors correspond to meaningful and uniquely determined latent factors and that these factors can be directly manipulated to achieve reliable control. However, classical results in latent variable modeling and causal inference show that such assumptions are often unjustified without additional structural constraints (Hyvärinen and Pajunen, 1999; Shimizu et al., 2006; Schölkopf et al., 2021; Locatello et al., 2019).
@@
Figure 1: Steering vector non-identifiability: The original vector $v$ and a perturbed vector $v+v_{\perp}$ with $v_{\perp}\in\mathrm{ker}(J)$ produce identical outputs since null-space components do not affect JvJv, the first-order change in output logits induced by steering.
We make three contributions. First, we provide a formal identifiability analysis of persona vector steering, proving that under white-box single-layer access, persona vectors are fundamentally non-identifiable: infinitely many geometrically distinct directions induce identical observable behavior via null-space ambiguity. Second, we empirically demonstrate that extracted steering vectors contain substantial non-identifiable components, with orthogonal perturbations achieving 95–100% of the original steering efficacy across multiple models and traits. Third, we validate through multi-environment experiments that this equivalence persists under distribution shift, confirming non-identifiability is a robust geometric property rather than just an artifact of limited prompt diversity. These results clarify when alignment methods can be trusted and which structural assumptions are required for reliable control.
@@ 2 Related Work
Causal and latent variable identifiability
Classical results show that latent variable models are non-identifiable without structural assumptions (Hyvärinen and Pajunen, 1999; Shimizu et al., 2006; Kruskal, 1977). Recent work in causal representation learning extends these ideas to deep learning settings (Schölkopf et al., 2021; Ahuja et al., 2022; Locatello et al., 2019), establishing conditions under which latent factors can be recovered from high-dimensional observations.
Probing and representation learning
Work on probing classifiers (Pimentel et al., 2020; Ravfogel et al., 2020; Elazar et al., 2021) shows that linear directions can succeed for reasons unrelated to target information. The linear representation hypothesis (Park et al., 2023; Elhage et al., 2022) suggests concepts correspond to directions in activation space but lacks identifiability guarantees.
Activation editing in LLMs
Methods such as representation engineering (Zou et al., 2023), contrastive activation addition (Rimsky et al., 2024; Turner et al., 2024) and activation patching (Meng et al., 2022) manipulate model internals but do not address whether control directions are uniquely determined. Work on neural network symmetries (Entezari et al., 2021; Dinh et al., 2017) motivates our null-space analysis.
Robustness and invariant representations
Work on distribution shift and domain adaptation (Rabanser et al., 2019; Lipton et al., 2018) establishes that learned representations often exploit spurious correlations. Methods for learning invariant representations (Arjovsky et al., 2019; Krueger et al., 2021) aim to extract features stable across domains. Recent work on multi-environment causal representation learning (Ahuja et al., 2022; Von Kügelgen et al., 2021; Rojas-Carulla et al., 2018) proves that diverse training environments enable identification of causal factors under structural assumptions.
Mechanistic interpretability
Work on mechanistic interpretability (Olsson et al., 2022; Elhage et al., 2021; Nanda et al., 2023) studies circuits and features in transformers, often assuming the existence of interpretable directions without formalizing identifiability constraints. Our theoretical analysis clarifies when such directions are well-defined.
@@ 3.1 Formal Model
Latent persona variable
@@ 3.2 Observational Regimes
  • Regime 1: Black-box input–output The researcher observes only $(x,y)$ pairs, where $y$ is generated text. There is no access to internal representations. This is the weakest regime and corresponds to behavioral evaluation.
  • Regime 2: White-box single-layer access The researcher can observe or manipulate activations $h_{\ell}(x)$ at a chosen layer $\ell$. This is the standard setting for most steering work and includes extracting vectors from contrastive prompt pairs.
@@ 3.4 Assumptions
We make the following assumptions explicit.
A1 (Smoothness) The functions $g_{\ell}$ and $F_{\ell\to L}$ are differentiable almost everywhere, enabling local linear approximation via Jacobians $J_{\ell}(x)=\frac{\partial o}{\partial h_{\ell}}(x)$.
A2 (Identifiable prompts) The prompt distribution $p(x)$ has sufficient variability to probe different aspects of the latent persona $z$. Formally, the support of $p(x)$ is rich enough that different persona values induce distinguishable activation patterns $h_{\ell}(x)$.
A3 (Non-degeneracy) The Jacobian $J_{\ell}(x)$ has rank at least $k\geq 1$ for typical $x\sim p(x)$, meaning steering can affect outputs. This excludes pathological cases where all perturbations to $h_{\ell}$ are ignored by subsequent layers.
We do not assume: (i) statistical independence between $z$ and $x$ (confounding is allowed), (ii) linearity of $g_{\ell}$ in $z$ or (iii) uniqueness of the latent representation without additional structure. These assumptions are mild and generally satisfied by standard transformer architectures under typical operating conditions.
@@ 4.1 Identifiability
Definition 1 (Parameter Identifiability)
@@
Definition 2 (Observational Equivalence)
Two steering vectors $v$ and $v^{\prime}$ are observationally equivalent in regime $\mathcal{R}$ if they produce identical distributions over all quantities observable in $\mathcal{R}$. For Regime 2 (white-box single-layer access):
@@
$\quad=F_{\ell\to L}(h_{\ell}(x)+\alpha v^{\prime})\;\forall x,\alpha.$
@@ 4.2 Symmetries and Gauge Freedom
Null space ambiguity
@@ 5 Main Results
Proposition 1
Under Assumptions A1–A3 within the local linear approximation, in Regime 2 (white-box single-layer access) without additional structural constraints, persona vectors are not identifiable. Specifically, for any steering vector $v\in\mathbb{R}^{d}$, there exist infinitely many vectors $v^{\prime}\not\propto v$ that are observationally equivalent under linearization.
Proof Sketch
We establish non-identifiability via the null-space ambiguity argument. A related weight-space symmetry result is given in (Appendix B.2).
Null-space ambiguity (primary mechanism). Under the local linear approximation (Equation 4), the steered output is
$o(x,v,\alpha)\approx o_{0}(x)+\alpha J_{\ell}(x)v.$ (7)
For any $v_{0}\in\ker(J_{\ell})$, i.e., $J_{\ell}v_{0}=0$, define $v^{\prime}=v+v_{0}$. Then
$o(x,v^{\prime},\alpha)$ $\approx o_{0}(x)+\alpha J_{\ell}(x)(v+v_{0})$ (8)
$=o_{0}(x)+\alpha J_{\ell}(x)v\approx o(x,v,\alpha),$
Since $\dim(\ker(J_{\ell}))=d-\mathrm{rank}(J_{\ell})$, any case where $\mathrm{rank}(J_{\ell})<d$ implies a non-trivial null space. In overparameterized language models the effective rank of $J_{\ell}$ is often smaller than $d$ (Oymak et al., 2019), yielding $\dim(\ker(J_{\ell}))\geq 1$ in practice. Hence infinitely many such $v_{0}$ exist, producing observationally equivalent vectors $v^{\prime}\not\propto v$.
Remark (Scope of linearization)
Proposition 1 is proved under the local linear approximation (Equation 4), which holds exactly for piecewise-linear activations (ReLU MLPs on activation-region interiors) and as a first-order approximation for smooth nonlinearities (GeLU, attention softmax). For small steering strengths $\alpha$, the nonlinear residual is $O(\alpha^{2})$ and empirically negligible. Extending the result to the exact nonlinear regime $o=F_{\ell\to L}(h_{\ell}+\alpha v)$ would require characterizing the Hessian structure of $F_{\ell\to L}$ and is left for future work.
Corollary 1.1 (Null-space equivalence)
Under the linear approximation $o\approx o_{0}+\alpha J_{\ell}v$, any vector $v^{\prime}=v+v_{0}$ where $v_{0}\in\ker(J_{\ell})$ is observationally equivalent to $v$ since $J_{\ell}v^{\prime}=J_{\ell}v+J_{\ell}v_{0}=J_{\ell}v$.
@@ 6 Empirical Validation
We now validate that the non-identifiability characterized in Proposition 1 (Section 5) manifests in contemporary language models. Our empirical experiments test a behavioral consequence of the theoretical prediction: the existence of large equivalence classes of semantically indistinguishable steering directions, without directly estimating the Jacobian null space. A complementary logit-level analysis in Appendix A further examines output distribution preservation beyond the semantic probes used in the main experiments.
@@ 6.1 Experimental Setup
Models and Layers
We evaluate two open-weight instruction-tuned language models to test generality across architectures and model scales: Qwen2.5-3B-Instruct (layer $\ell=12$) and Llama-3.1-8B-Instruct (layer $\ell=16$) (Yang et al., 2024; Meta AI, 2024). For both models, we focus on middle layers ($\ell=L/2$), following standard practice (Chen et al., 2025; Konen et al., 2024; Sun et al., 2025) in steering research.
Persona traits
@@
Steering vector extraction
@@
Semantic probes
We evaluate semantic equivalence using trait-specific scoring functions $\phi(o)$ that map generated text to real-valued scores. For formality, politeness and humor, we use lexical heuristics based on formal/informal markers, polite/rude markers and humorous/serious markers respectively, combined with sentence length and stylistic features. These return scores in $[0,1]$ where 1 is maximally formal/polite/humorous. Direct distributional evidence is provided by the logit-level analysis in Appendix A.
Orthogonality test methodology
@@
Sample size design
@@
Scale invariance test
@@ 6.2 Orthogonal Perturbation Test Results
Table 1: Empirical validation of non-identifiability across models, traits, and sample sizes. Cohen’s $d$ measures the effect size between $v$ and $v+v_{\perp}$ (smaller = more equivalent). All results show negligible differences ($d<0.2$), supporting observational equivalence.
@@
With $n=10$ seeds, the mean Cohen’s $d$ is 0.080 for Qwen2.5-3B and 0.100 for Llama-3.1-8B, both well below the threshold for small effects ($d=0.2$) and firmly in the negligible range. Near-zero correlations (e.g., $0.044$, 0.032-0.032) reflect probe ceiling effects rather than independence; Cohen’s $d$ is the more reliable metric in these cases. The “Perp-Only Effect” column measures the efficacy of steering with pure orthogonal components ($v_{\perp}$ alone, without $v$) relative to the extracted vector—values near $100\%$ indicate that orthogonal components achieve equivalent behavioral impact. Values marginally above $100\%$ reflect probe ceiling effects at the upper boundary of the scoring range.
Statistical stability across sample sizes The convergence between $n=5$ and $n=10$ results demonstrates robustness across both models. For Qwen2.5-3B, effect sizes change by <0.07<0.07 as sample size increases, with all traits showing tightening confidence intervals: formality ($0.144\to 0.075$), politeness ($0.112\to 0.092$), humor ($0.103\to 0.072$). For Llama-3.1-8B, we observe similar stability: formality ($0.052\to 0.096$), politeness ($0.074\to 0.085$), humor ($0.159\to 0.119$). This consistency demonstrates that the observed equivalence is not a sampling artifact but a stable property of the steering geometry.
Cross-model consistency The close agreement between Qwen2.5-3B ($d=0.080$) and Llama-3.1-8B ($d=0.100$) demonstrates that observational equivalence is not model-specific. Despite differences in architecture, scale (3B vs. 8B parameters) and hidden dimension ($d=2048$ vs. $d=4096$), both models exhibit nearly identical patterns of non-identifiability. The consistency across traits within each model further confirms that the phenomenon is general rather than an artifact of specific semantic domains or model implementations.
Visualizing orthogonal component equivalence Figure 2 illustrates the perp-only effect ratios across all traits and models using $n=10$ orthogonal seeds. The tight clustering around the perfect equivalence line (dashed red at $1.0$) demonstrates that pure orthogonal components achieve nearly identical steering efficacy to the extracted vectors.
@@
Figure 2: Perp-only effect ratios ($v_{\perp}$ efficacy / $v$ efficacy) for $n=10$ orthogonal seeds. Values near 1.0 indicate equivalence.